(1 point) Find the solution of y + 14y +48 y = 24 e-St with y(0) = 3 and y0) = 4. y =
Answer #1

Given y + 18 Y +48 Y = QH é o consider y+ lh yt 489 = 2 test apply Laplace Transfeom on both sides d{ + (Hy + 84} = tea- 24 + t B St6 - st5 - Sto - = 8 --IG (5+5) (5+6)(548) A = BE24.Com -Ica) Categy -- YCS) = St8 =H + 14 -+ + S+6 S45 St6 S8 YC